Circular No. 6393

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1655–40 R. Remillard, H. Bradt, W. Cui, A. Levine, E. Morgan, B. Shirey, and

D. Smith, Massachusetts Institute of Technology, report: “Observations with the RXTE All Sky Monitor (ASM) reveal a recent recurrence of bright x-ray emission from GRO J1655–40, a dynamical blackhole binary (Bailyn et al. 1995, Nature 378, 157) that began x-ray eruptions during 1994 and became a source of relativistic radio jets (Tingay et al. 1995, Nature 374, 141). Combining the ASM measurements at 2–12 keV in 1-day time bins (UT), we see an upper limit of 12 mCrab (3σ) on 1996 Apr. 23, followed by detections of 71 ± 15 mCrab on Apr. 25, 420 ± 30 mCrab on Apr. 26, 586 ± 15 mCrab on Apr. 27, 802 ± 9 mCrab on Apr. 28, and 1077 ± 30 mCrab on Apr. 29. Between Apr. 30 and May 5, the x-ray flux ranges between 1.3 and 1.5 Crab. Observations at all wavelengths are strongly encouraged.”

COMET C/1996 B2 (HYAKUTAKE) F. E. Marshall, C. Lisse, M. Mumma, and R. Petre, NASA Goddard

Space Flight Center; and K. Dennerl, J. Englhauser, and J. Trümper, Max-Planck-Institut für Extraterrestrische Physik, Garching, report: “The xray emission from comet C/1996 B2 is primarily at energies < 2 keV. No evidence for emission in the band 2–10 keV was seen during a fortuitous scan lasting 24 s across the comet on Mar. 31.2836 UT with the Proportional Counter Array onboard the Rossi X-ray Timing Explorer. The 90-percent-confidence upper limit for the flux in the band 2–10 keV, assuming a powerlaw spectrum with photon index 2.0, is 1.6 × 10^{−11} erg s^{−1} cm^{−2}. The limit is 1.7 × 10^{−11} erg s^{−1} cm^{−2} for an index of 3.0. During this observation, the comet was 0.22 AU from the earth and 0.91 AU from the sun. Assuming the comet’s luminosity in the band 0.1–2 keV was the same as during the nonflaring intervals of the ROSAT observation on Mar. 26–28 (IAUC 6373), we can rule out spectral shapes harder than a power law with photon index 2.0 or thermal bremsstrahlung emission with a temperature of 2.5 keV. For both these assumed spectral shapes, > 50 percent of the 0.1- to 10-keV luminosity is at energies < 1.25 keV.”

FG SAGITTAE S. Shugarov, Sternberg Astronomical Institute, Moscow, reports that

this variable has faded to mag V ∼ 16 and B ∼ 17 on Apr. 24.04 UT. A CCD image by R. Royer, Lakewood, CA, yields mag 16–16.5 on May 7.36.

1996 May 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green